A spring can be stretched a distance of 30 cm with an applied force of 1 N. If the same spring is stretched same distance with 1 N force, how much energy is stored in the spring?
Answer #1

For spring

F = k x

1 N = (k) × (30 cm)

1 N = (k) × (0.3 m)

k = (1 N)/(0.3 m) = 3.33 N/m

Energy stored in spring is given by,

U = (1/2) k x^2

= (1/2) × 3.33 × (0.3)2

=  0.15 J